OpenStudy (anonymous):

You need to drill a hole in a piece of metal for a machine shop project. The specifications indicate that the diameter of the hole should be at least 0.355 inch and at most 0.375 inch. Write an absolute value inequality that represents the acceptable diameters of the hole.

OpenStudy (helder_edwin):

\[ \large 0.355\leq D\leq0.375 \]

OpenStudy (anonymous):

ok for the answer it said that it is [x- 0.365]≤ 0.010 * []=ablsolute value How did they get that?

OpenStudy (helder_edwin):

sorry i had to for a while: \[ \large 0.355\leq D\leq0.375 \] substracting 0.365 u get \[ \large -0.010\leq D-0.365\leq0.010 \] from this \[ \large |D-0.365|\leq0.010 \]

OpenStudy (anonymous):

how did you know to subtract 0.365?

OpenStudy (helder_edwin):

that's the only number with which u get the same number (but with different sign) on both extremes.

OpenStudy (anonymous):

so what is the equation to find 0.365?

OpenStudy (helder_edwin):

what is the middle point between 0.355 and 0.375

OpenStudy (helder_edwin):

??

OpenStudy (anonymous):

oh so you add them both and divide by 2 right so it would = to 0.365

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

oh ok and how did you isolate the -0.10 & 0.10 to be at the end of the inequality equation?

OpenStudy (anonymous):

do you multiply it them?

OpenStudy (anonymous):

*do you multiply them ? srry

OpenStudy (helder_edwin):

no i substracted \[ \large 0.375-0.365=0.010 \] and \[ \large 0.355-0.365=-0.010 \]

OpenStudy (anonymous):

yea but the final equation is \[| D- 0.365|\le 0.010\] , howd you get from -0.10<=D-0.365<=0.10?

OpenStudy (helder_edwin):

oh. it is a theorem from calculus (absolute value): \[ \large |a|\leq b\Leftrightarrow\begin{cases} b\geq0\\ \text{and}\\ -b\leq a\leq b \end{cases} \]

OpenStudy (anonymous):

oh well i haven't learned that yet, that's why i didn't know

OpenStudy (helder_edwin):

well. now u know :)

OpenStudy (anonymous):

Well i know how it looks like but, not how it works .

OpenStudy (anonymous):

Ill just wait till class so the teacher can explain because it looks to complicated for you to explain through here. So thanks for the help!

OpenStudy (helder_edwin):

u r welcome